OpenStudy (anonymous):

Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. Point pairs (5, 1), (–3, 4) (0, –2), (3, 2) (–2, –1), (1, 2)

hero (hero):

Use the slope formula to find the slope, m, between two points: \[m = \frac{y_2 - y_1}{x_2 - x_1}\] Then use the point slope formula \[y - y_1 = m(x-x_1)\] to input the slope, m, and the coordinates of the first point into that formula.Then, simplify in the form y = mx + b.